The length of a rectangle is 6 feet more than the width of the rectangle. The length is 32 feet. what is the width?
Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52781)

.

Let w be the width.


Then you have this equation


    w + (w+6) = 32/2,   or


    w + (w+6) = 16.


At this point, I step away and leave you to complete the solution on your own.

You may inform me about your progress.



Answer by MathTherapy(10552)


The length of a rectangle is 6 feet more than the width of the rectangle. The length is 32 feet. what is the width?
Length: W + 6, with W being the width
Length: 32 feet
You're telling me that you can't find the width based on the given info.?